The bisector of the angle A at the base of the trapezoid ABCD is perpendicular to it's diagonal and crosses the side CD,at the point E and CE : ED=2:3,Find the ratio of BC:CK where k is the intersection point of the bisector with line BC.
1
Expert's answer
2020-08-11T16:31:46-0400



Consider a trapezoid ABCD shown in the diagram above.


Let AE be the bisector of "\angle A" where "\angle DAE = x" so "\angle BAK = x" as per the angular bisector property.


As per the question AE is perpendicular to the diagonal BD

"\therefore" "\angle ADB = \angle ABD = 90\degree - x"


"\angle KCE = x" as AE is intersecting two parallel lines AB and DE so incident angles are equal.


Let "\angle CKE = a" so "\angle BKA = a" as vertically opposite angles are equal.


"\therefore" we can say that "\triangle ABK \cong \triangle CEK" by AAA congruent property.


"\therefore" CE/AB = CK/BK


Now we can write BK as (BC-CK).


Also AB= AD as "\triangle ADB" is an isosceles triangle.


Also "\triangle ADE" is an isosceles triangle because two base angles are equal to "x"


so AD = ED which means AB = ED


Now coming back to below equation and putting the values we get.



CE/AB = CK/BK


CE/ED = CK/(BC-CK)


ED/CE = (BC-CK)/CK


ED/CE = (BC/CK) - 1


From the question ED/CE = 3/2


"\therefore" BC/CK = (3/2) + 1

= 5/2